OpenStudy (anonymous):

Multiply and Simplify: 2√5(3√2 + 3√5)

OpenStudy (adamaero):

\[x \sqrt{y}(z \sqrt{q}) = xz \sqrt{yq}\] an example of foil: http://cnx.org/resources/52ce8781a2c06cd1114e711ad226e447/C08_S8-5_P27_003.jpg

OpenStudy (anonymous):

\[\sqrt{y}\times \sqrt{q}=\sqrt{yq},only ~if~both~are~not~ together~negative.\]

OpenStudy (anonymous):

What would the solution be